F-2 was successfully checked on January 26th after an intevention that lasted for most of January. New JH and HK gratings were installed and its performance will be assessed in the following days. Additionally, the J_low filter is now permanently installed.
An intervention of Flamingos-2 is planned between January 3rd and 21st where the instrument will be offline. The following tasks are planned:
MOS wheel check and preventive work.
Replace the JH/HK spectroscopic filters with higher throughput ones.
Modify the Lyot wheel to allow the installation of the J-low and Y filters instead of the Hartmann masks.
Increase the number of layers in the dewars insulation, to allow more thermal stability.
After a thermal cycle performed over the weekend, the detector amplifier #5 is back to its normal perfromance. CCD1 CTE is also nominal, and OIWFS performance is OK after the successful repair that had been performed during September. Therefore GMOS-S is back at full capacity.
After a successful realignment of its optics, the first Zorro run of the semester happened between Sept 16 and 22. Zorro has both cameras operative and waiting for proposals for 22A!
GMOS was installed on the telescope, but it is not available yet due to the issue with CCD2 on the GMOS detector array, which features enhanced noise structure. The troubleshooting will be carried away during this week.
Zorro will be removed on Sept 1st for its optics realignment between Sept 8 to 10th. It will be installed back on the telescope between Sept 13 and 15th, and be ready before the Zorro run starting on the 16th.
UPDATE (Sep 03): the repair work will extend through mid-September. The reason for this is that one of the optical components of the OI probe needs to be re-attached to the pickoff arm after it was found loose. The work will also involve the corresponding optical alignment check.
Due to a mechanical failure, The GMOS-S OIWFS will be out of service at least for the rest of August. A repair plan at the CPO Instrument Lab is being worked on. As a consequence of this situation, GMOS-S will be available only with PWFS2 as guiding option.
The Zorro blue filter wheel, that had been faulty since last March, was successfully replaced last week. An unfortunate side-effect of the procedure is that the red and blue cameras are now misaligned by about 6 arcsecods, a shift bigger than the speckle field-of-view, which makes simultaneous observations with both cameras in speckle mode impossible. Both cameras are working fine, only independently. New PIs should therefore indicate which camera they want to use. Both cameras can still be used, but this will imply doubling the requested time.
It was recently reported that the observed counts in spectra seemed to be lower than expected on CCD1 for some programs - also, odd emission-like features (‘spikes’) were noticed in the CCD1 portion on stellar spectra that are known to be featureless. All this applies only to data acquired during the ‘bad-CTE’ periods of CCD1, i.e. between:
NIFS has a cryogenic issue which started at the time of the earthquake and power outage on Monday 5th July. This problem has developed and it is now clear that the instrument must be warmed up completely to effect mitigation. NIFS will remain off the telescope until this is completed; NIRI will remain on for the time being.